Suppose you do a carry trade for one semester in USDMXN in which you make a short sale of 1,000,000 USD at 22.10 MXN/USD. If you closed this trade at 22.05 MXN/USD and the effective 3-month rates for loans and deposits is 1.2500% for MXN and 0.0625% for USD, what are your profits? (Use six decimals for your calculations but round your final answer to the nearest unit of dollar)

A) 26,230.0000

B) 262,306.0000

C) 56,810.0000

D) -26,171.0000
Answer #1

Option A

Proceeds on sale of 1,000,000 USD=1000000*22.10= MXN 22100000

Return on investment=1.25% for 3 months

Value after the semester=Investment value*(1+interest for 3 months)^2=22100000*(1+0.0125)^2=MXN 2265953.125

$ proceeds when trade is closed=2265953.125/22.05=$1027480.86734

$ payable after 3 months when $ borrowed=1000000*(1+0.000625)^2=$1001250.39062

Profit=$ from trade-$ payable=$1027480.86734-1001250.39062=$26230.47672=$26230 (approx.)